OpenStudy (anonymous):

ok so i am getting paid $9.14 an hour i work 6 hours a day for 4 days a week if i get paid every two weeks how much would be my first paycheck and i get paid every 2 weeks

OpenStudy (jdoe0001):

well... how much do you get paid for 1 week then? $9.14 per hour you work 6hrs a day so $9.14 * 6 in 1 day you work 4 days a week, so in 1 day is $9.14 * 6 how about in 4 days?

OpenStudy (jdoe0001):

that'd give you how much in 1 week since you get paid every 2weeks, then your pay would be the pay for 2 weeks :)

OpenStudy (jdoe0001):

$9.14 is not horrible btw... though close to minimum wage these days not really great per se

OpenStudy (anonymous):

so wat will i be getting paid every 2 weeks

OpenStudy (jdoe0001):

well... what did you get for 1 week?

OpenStudy (anonymous):

idk ik im getting paid 9.14 an hour 6 hours a day for 4 days a week

OpenStudy (jdoe0001):

yes.... so 9.14 for 1 hr so for 6 hours that'd \(\begin{array}{llll} 9.14\\9.14\\9.14\\9.14\\9.14\\9.14\\ \hline\\ \square ? \end{array}\implies 9.14\times 6\)

OpenStudy (anonymous):

thats 54.84 a day ok now wat

OpenStudy (jdoe0001):

so ok... so for 1 day is $54.84 so for 4 days that'd be \(\begin{array}{rllll} 54.84\\54.84\\54.84\\+54.84 \\\hline\\ \square ? \end{array}\implies 54.84\times 4\)

OpenStudy (anonymous):

219.36 now wat

OpenStudy (jdoe0001):

yeap... that's 4 days of work, at 6hrs each day at $9.14 per hour that's 1 week you get paid every two weeks so your cheque would be twice that much :)

OpenStudy (anonymous):

so wat i needa do to find out every 2 weeks

OpenStudy (jdoe0001):

well \(\begin{array}{rllll} 219.36\\219.36 \\\hline\\ \square ? \end{array}\implies 219.36\times 2\)

OpenStudy (anonymous):

438.72

OpenStudy (anonymous):

i also worked 6 extra hours so do i Times 16

OpenStudy (jdoe0001):

hmmm 6 extra hours?

OpenStudy (anonymous):

no 16 extra hours so 438.72 times 16 right

OpenStudy (jdoe0001):

well... each hour is 9.14 so 6 extra hours will be \(\begin{array}{rllll} 9.14\\9.14\\9.14\\9.14\\9.14\\+9.14\\ \hline\\ \square ? \end{array}\implies 9.14\times 6\)

OpenStudy (jdoe0001):

hmmm 16 extra hours? besides the 2weeks worked already?

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

well every hour is 9.14 so 16 hours will be 16 * 9.14 so you'd add that to the 438.72

OpenStudy (anonymous):

so the answer is 584.96

OpenStudy (jdoe0001):

yeap